OpenStudy (anonymous):

6 5/8 divided by 6 1/4 . . . ? (:

OpenStudy (anonymous):

53/50 because you need to do the reciprocal of the other. 53/8 x 4/25

OpenStudy (amistre64):

\[\frac{6+\frac{5}{8}}{25/4}\] \[\frac{24+\frac{5}{2}}{25}\] \[\frac{24}{25}+\frac{\frac{5}{2}}{25}\] \[\frac{24}{25}+\frac{\frac{1}{2}}{5}\] \[\frac{24}{25}+\frac{1}{10}\] maybe